solve by any appropriate method. if it is not possible to obtain one solution, state the reason. write the solution in the form (x,y).
1 + x - y = y + 4
4(x-y) = 3 - x

Answer by stanbon(75887)

solve by any appropriate method. if it is not possible to obtain one solution, state the reason. write the solution in the form (x,y).
1 + x - y = y + 4
4(x-y) = 3 - x
-------
Rearrange:
2y - x = -3
4y -5x = -3
------
Modify:
4y - 2x = -6
4y - 5x = -3
-----
Subtract and solve for "x":
3x = -3
x = -1
-------------
Solve for "y":
2y -x = -3
2y + 1 = -3
y = -2
=====
Solution: (-1,-2)
=====================
Cheers,
Stan H.
=====================